OpenStudy (anonymous):

Positive integers B and C satisfy B(B-c)=23. What is the value of C? show all work

OpenStudy (anonymous):

You cannot solve for C with just this information.

OpenStudy (anonymous):

Positive integers B and C satisfy B(B-C)=23. What is the value of C? can you solve it now?

OpenStudy (anonymous):

No. There are many solutions to the equation \[ C = \frac{B^2-23}{B}\] The only restriction is that \[ B^2-23 > 0 \rightarrow B > \sqrt{23} \] Any value for B greater than the square root of 23 will yield a different valid value for C.

OpenStudy (nikvist):

\[B=23,C=22\]